OpenStudy (anonymous):

2x/x^2+1=2/x+2 what is the value of x?

OpenStudy (campbell_st):

so is the question \[\frac{2x}{x^2 + 1} = \frac{2}{x + 2}\]

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

The solution is x =1/2

OpenStudy (campbell_st):

ok cross multiply to eliminate the denominators and its \[2x(x + 2) = 2(x^2 + 1)\] if you distribute you can then start to solve for x hope it makes sense

OpenStudy (campbell_st):

there you go... someone always gives the answer

OpenStudy (anonymous):

Okay, thank you both so much!